Summary of the 3mm GMVA Session May 6-11, 2010 --------------------------------------------------- In this session the following stations participated: Eb, On, Mh, PV, PB, 8xVLBA An ftp-fringe test performed on May 7 from 05:17:00 UT to 05:17:20 UT on the Ef, PB, PV baselines gave excellent fringes on 3C454.3 in both polarisations. SNRs were 700-800 on EB-PB and EB-PV and SNR 2500 on PV-PB. Note: On April 29, a 86 GHz fringe test was performed between On-Mh and Yb (Yebes) on 3C454.3, which has about 30 Jy. Mh has a new MMIC based 86 GHz receiver. At this early stage, the new Yb antenna and receiver are not yet fully optimized, and the antenna gain is still low. The feed polarisation at Yb was RCP, however with large cross-terms (to be fixed). The following SNR rations were achieved: On-Yb 110, On-Mh 50, Mh-Yb 51-68. The session was affected by bad to moderate weather at Effelsberg, with rain on May 6, overcast and cloudy conditions on May 7-10 and rain during the last 12 hrs on May 11. The weather at IRAM was better, with some downtime because of wind (PV) and snow (PdB). The weather conditions at the VLBA were mostly good, except for NL, which had some rain. More detailed log information: ------------------------------ Effelsberg: May 6 22 UT rain (1502+102) May 7 12 UT overcast, pointing ok (NRAO150) May 8 0 UT partly cloudy, pointing ok (CygA) 12 UT overcast, pointing problems (CygA) 13 UT cloudy, pointing ok (OJ287) 20 UT clear sky May 9 8 UT overcast (3C345, 3C446), Tsys ~180 K 16 UT overcast (3C111), pointing ok May 10 6 UT overcast, pointing ok (BLLAC) 18 UT light clouds (0420-014, 0528+135) May 11 5 UT cloudy, pointing ok (0716+714) 16 UT heavy rain, bad data (0420-014, 0528+135) Pico Veleta: May 6 22 UT cloudy but ok May 8 2 UT ok 19 UT wind stop (OJ287, 3C111) May 9 8 UT wind stop (3C111) 19 UT back observing (3C273/279), Tsys 125 K May 10 6 UT ok (BLLAC) 11 UT bad weather (BLLAC) May 11 3 UT 2h45-3h45 MK5 problem, bad disk 5 UT snow stop (3C454.3, 0528+134, 0420-014) Plateau de Bure: May 6 22 UT ok May 7 19 UT bad weather (CygA) May 8 2 UT PdB resumes, weather improved 13 UT snow stop May 9 9 UT bad weather, no phasecal in RCP 16 UT source is in sun avoidance May 10 6 UT ok (BLLAC) May 11 0 UT ok 9 UT rain stop Onsala: all projects observed, no known major technical problems. Some rain, particularly on May 8. Metsahovi: report from Ari Mujunen: Rain on day 127. All time quite cloudy, relative humidity is constantly >90%. During scan No0163 (127.19:40) the Mk4 formatter spontaneously reset itself, including its clock. A power glitch would otherwise be a probable reason but other devices / computers in the same UPS system did not reset, so the reason remains a mystery until rack power supplies and UPS devices can be manipulated and tested at will. The formatter clock was reset later (128.08:17) but that helped very little, as it will become evident in the following. During scan No0180 (127.22:45) the Mark5A started to "throttle" recording. This went unnoticed since this condition doesn't emit a real FS error message labeled with "ERROR" in the experiment log. Also the recording did not stop completely, the record byte pointer continued to increase, just too slowly. Apparently Mark5A didn't like the new 6TB disk pack built with Samsung 750GB disks---it passed the SSErase but when the pack was removed it seemed surprisingly hot although the pack fan was working just fine. This is of course just guessing why the Mark5A doesn't like this pack. The last scan recorded with "throttle" was No0383---during this the FS stopped running the schedule and also stopped responding to manual oprin commands ('status', 'terminate'). It didn't crash, though, and thus the windows on screen looked just normal. The reason for this is not known, though I'm suspecting that a throttling Mark5A program maybe becomes unresponsive to the FS and may slow it down / halt it, maybe? Anyway, a complete restart was performed today (130.07:30) and we are continuing with scan No0586 normally, though the weather is really cloudy and on the verge of starting to rain. It remains to be seen if the (supposedly) okay 1TB of recording in the beginning of the 6TB SATA pack can be salvaged. Obviously I don't want to perform disk pack tests while the tail of C101A is still running since these tests with potentially faulty drives/packs frequently crash the Mark5A completely. The worst case would be if the pack refuses to mount at all---a slightly better scenario would be that the pack will work for a limited time before it overheats itself, we'll see. VLBA: weather May 6 0 UT all good weather May 7 15 UT NL rain, other ok May 8 0 UT NL cloudy, other ok 15 UT NL, FD, PT cloudy 20 UT NL, FD, PT cloudy others clear May 9 8 UT NL light clouds 20 UT all ok May 10 18 UT all ok, NL partly cloudy May 11 16 UT all ok, NL clouds more detailed log information: VERY LONG BASELINE ARRAY OBSERVING LOG -------------------------------------- Project: C101A Observer: Krichbaum, T. Project type: GEO Obs filename: c101acrd.* Date/Day: 2010MAY06/126 Ants Scheduled: NL FD LA PT KP OV BR MK =UT-Time===Comment===============================================MF#===%AD==AMD= Operator is Heidi Medlin GPS delta, UT1-UTC(-0.03), vclock ok %LA 3mm will have high Tsys but data is okay. 22676 2357 Begin 2357-1853 %BR 7mm LCP system temperatures are roughly 26577 25 1724 two times higher than RCP system temperatures. 2357-1853 %FD 3mm cryo is a little warm, usually below 30K but sometimes spiking just a little above. 0000 Date is 2010MAY07/127. 0104-0107 *PT power glitch COM 100 3 0112 *PT 35+ pmh wind caution 0143-0313 *PT power glitch COM 100 90 0545 New operator: T. Perreault 0836-0919 %NL wea: raining 1027-1244 %NL wea: raining 1103-1210 &FD unclearable EL2 armature/circuit breaker fault; 26827 100 67 site tech called out. Antenna placed in STANDBY until techs arrive. 1153 FD tech on site, has taken the antenna 1237-1245 *FD Mark5 crashed when new pack was inserted into REC 100 8 bank A. 1342 New Operator: Paul Padilla 1923-1938 *PT stowed, site tech checking vertex room temps VAC 100 15 1946 %PT A Rack temps slightly high due to warm vertex room 1956-2017 *PT stowed, site tech opening another door to VAC 100 21 try to keep vertex room temps under control 2138 New operator is Heidi Medlin 2314-2320 *MK rotation brake fault FRM 100 6 0000 Date is 2010MAY08/128. 0105-0107 *KP power glitch COM 100 2 0543 New operator: T. Perreault 1343 Operator is A. Sowinski. 0000 Date is 2010MAY09/129. 0144 New operator is Betty Ragan 1346 Operator is A. Sowinski. 2031 %OV Windy. 2332 *FD Raining. 0000 Date is 2010MAY10/130. 0140 New operator is Betty Ragan. 1045 %PT windy. 1054-1056 *PT autostowed. Power glitch. COM 100 2 1127 %LA windy. 1346 Operator is A. Sowinski. 1354 LA wea: broken low clouds, dry. 1425 BR wea: high thin cloud cover, dry. 1438 PT wea: clear, windy, dry. 1455 OV wea: partly cloudy, dry. 1521 %FD Windy. 1532 LA wea: clear, dry. 1619-2018 *PT Antenna stowed for work on vertex room A/C. VAC 100 239 1808-1815 *FD Windstow. Will repoint in a few minutes. WEA 100 7 1815 KP wea: clear, dry. 1924 MK wea: clear. 1924 *NL Raining. 1959-2007 *FD Windstow again. WEA 100 8 2155 New Operator is Alan Kerr 0000 Date is 2010MAY11/131. 0029-0147 %FD EL #2 armature/electroics CB (again). site 26838 tech called out. don't know if data is affected. the ACU is usually lagging by less than 6 seconds-of-arc. 0140-0147 *FD out of point. site tech working on EL CB. 26838 100 7 0553 New operator is Heidi Medlin 0558 %OV 35+ mph wind caution 0630-0631 *PT HC1 FIFO error 21522 100 1 0845 %PT 35+ mph wind caution 1025-1406 *KP lost power to A rack causing problems with the 26840 100 221 B rack, SYNs, and BBCs. called out site tech 1207 %OV 35+ mph wind caution 1333-1406 KP stowed for site techs to work on antenna 26840 1347 Operator is A. Sowinski. 1348 LA wea: clear, dry. 1401 BR wea: clear, dry. 1409 PT wea: clear, light wind, dry. 1503 OV wea: cloudy, dry. 1734-1815 *BR Power failure at site. Antenna stowed. COM 100 41 1750-1751 *PT Autostow, repointed. ACU 100 1 1842-1843 *PT Autostow, repointed. COM 100 1 1853 End. Downtime Summary: Total downtime : 2464 min Percentage downtime of observing: 4.5% Average downtime per hour : 2.7 min Total scheduled observing time (# Antennas): 55168 min (8) Notes: * = Entries where data was affected. % = Entries where data may have been affected. & = Entries where the site tech was called out. WEA = Weather entries. MF# = Maintenance form or major downtime category associated with a problem. %AD = The percentage of an antenna affected by a problem. AMD = Total antenna-minutes downtime for a problem. Tsys = System Temperature (TP/SP x Tcal/2) ACU = Antenna Control Unit FRM = Focus/Rotation Mount RFI = Radio Frequency Interference VME = Site control computer CB = Circuit Breaker vclock = Program that compares site clock time to a standard.